For a solution with pH = 9.45, the [OH-] is:

A. 1.0 x 10-7
B. 3.5 x 10-10
C. 2.8 x 10-5

The answer is C. 2.8x10^-5

How do you find this answer?
Answer #1

[H307] = 10th -9.45 [Hoof) = 109.45 = 3.54813 39x10-10 m (H3 ot] x [ot] = 10-24 at 256 [ot] = loty [H₂ot] [oH)= loty 3-548133